Project

General

Profile

Task #24754

Remove & Repair HWR Pressure loop Signal conditioning chassis

Added by John Dusatko 4 months ago. Updated 4 months ago.

Status:
Resolved
Priority:
Urgent
Assignee:
Start date:
08/11/2020
Due date:
% Done:

100%

Estimated time:
Spent time:
Duration:

Description

Cavity 1 pressure loop controls were commissioned on 8/3 and shown to work correctly. On 8/7, further testing began of the pressure/resonance control loop in frequency mode revealed that the pressure controls were no longer working correctly: When the loop was placed in manual (open loop) mode, the valve controls were set to -32767 and +32767, which should have produced a pressure transducer readback of +3.3V and +0.7V respectively. Instead, the readings showed: +3.3V (OK) and +2.3V (not ok). Further investigation by removing the Burndy cable and probing the connector pins on the chassis showed that the pressure valve control voltage measured -4.28V, when instead it should have measured closer to +2.5V and never be negative. Suspect that the U1 amplifier for that channel is damaged. Recommend replacement of chassis or PCB in chassis.

History

#1 Updated by John Dusatko 4 months ago

  • Tracker changed from Bug to Task

#2 Updated by Brian Chase 4 months ago

It would be a good idea to look at the valve end of the cable to make sure that something isn't shorting and backfeeding the controller. We currently have no mechanism for failure. it should a high z input.

#3 Updated by Daniel Klepec 4 months ago

  • % Done changed from 0 to 50

Replaced U1A, LT1995, which is associated with vacuum valve drive 1. John is in the process of checking it out.

#4 Updated by Daniel Klepec 4 months ago

  • % Done changed from 50 to 100
  • Status changed from New to Resolved

All channels are working per John including Channel 1 which was broke.

Also available in: Atom PDF